OpenStudy (anonymous):

0.02(y-1)+0.12y=0.16y-0.01(10)

OpenStudy (anonymous):

multiply by 50 (y-1)+6y=8y-5 4=y

OpenStudy (anonymous):

just need to simplify it 0.02y-0.02+0.12y=0.16y-0.1 0.14y-0.02=0.16y-0.1 -0.02y=-0.08 y=0.08/0.02=4